Find w=a +bi = VE , where a and b are real numbers. 7601 - 11 Answer: w= w=
Answer #1

J6ot-U : a + b = 10 Squaring on both sides (vooi-ll). (a-tbi) ? Goi-1 . at(bij + 2 abi 60i-11 (a-b)+zabi Comparing both sid0 a(a + 3). 05 ( ) (136) la? 25) = 0 I a +36=0 a² 25.0 a = -36 , a²= 25 a: 5-36 , a= Vis a=tbi, a=ts O for at bé , b: 20 -51